OpenStudy (muhammadturawa):

Find dy/dx for the following: lny=y-x^2

OpenStudy (anonymous):

do you know implicit derivatives?

OpenStudy (muhammadturawa):

Yes, I end up with this:

OpenStudy (muhammadturawa):

(1/y)(dy/dx) = (dy/dx)-2x

OpenStudy (anonymous):

yeah then subtract dy/dx from the right side and factor (1/y)(dy/dx) - (dy/dx) = -2x (dy/dx)(1/y - 1) = -2x

OpenStudy (anonymous):

dy/dx= - 2x/(1/y - 1)